### v2

# Spectrally Deconfounded Random Forests

Abstract
We introduce a modification of Random Forests to estimate functions when unobserved confounding variables are present. The technique is tailored for high-dimensional settings with many observed covariates. We employ spectral deconfounding techniques to minimize a deconfounded version of the least squares objective, resulting in the Spectrally Deconfounded Random Forests (SDForests). We show how the omitted variable bias in estimating a direct effect goes to zero, assuming dense confounding and high-dimensional data. We compare the performance of SDForests with that of classical Random Forests in a simulation study and a semi-synthetic setting using single-cell gene expression data. Empirical results suggest that SDForests outperform classical Random Forests in estimating the direct regression function, even if the theoretical assumptions are not perfectly met, and that SDForests and classical Random Forests have comparable performance in the non-confounded case. We provide an R-Package for SDForest, and supplementary materials for this article are available online.
Keywords: Causal Inference, Confounding, High-dimensional setting, Omitted variable bias, Regression
1 Introduction
Random Forests (Breiman 2001) and their variations, such as Random Survival Forests (Hothorn 2005, Taylor 2011), Quantile Regression Forests (Meinshausen 2006), or distributional versions of Random Forests (Hothorn & Zeileis 2021, Ćevid et al. 2022) are successfully applied to a wide range of datasets. In many cases of observational data, however, problems with “omitted variable bias” (Cinelli & Hazlett 2020, Wilms et al. 2021) occur. This means that a bias is induced in estimating relationships using standard Random Forest versions when covariates that correlate with other covariates and the response are not observed and included. In the setting of causality, this can be viewed as a confounded causal relationship with unobserved confounders (Pearl 2009, Peters et al. 2016). A popular approach to deal with unobserved confounding is to use instrumental variables (IV) regression techniques (Bowden et al. 1990, Angrist et al. 1996, Stock & Trebbi 2003). Finding enough strong and valid instrumental variables can be challenging, especially if many covariates with potential effects on the response are observed since the number of instruments must be as large as the number of effective covariates. Another possible way of reducing the hidden confounding bias, which we will adopt here, is to make some kind of “dense confounding effect” assumption, meaning that the non-observed factors or confounders affect most of the covariates. Then, a standard approach is to estimate the hidden confounding using methods from high-dimensional factor analysis (Bai 2003) and explicitly adjust for them, see for example Leek & Storey (2007), Gagnon-Bartsch & Speed (2012), Fan et al. (2024) for approaches in this direction. Instead of estimating the latent factors explicitly, one can adjust for them implicitly using spectral transformations (Ćevid et al. 2020). Applying such spectral transformations, especially the trim transform, which we introduce later, does not require any tuning and is computationally very fast since it is a simple and explicit function of the singular value decomposition of the design matrix.
1.1 Our Contribution
Our contribution is a Random Forest algorithm that is able to address, at least partially, the problem of hidden confounding. Our proposal combines the great advantage and flexibility of standard Random Forests with spectral deconfounding techniques for dealing with the bias from unobserved factors or confounding. The latter was originally proposed for linear models by Ćevid et al. (2020), Guo et al. (2022) and was further developed for nonlinear additive models by Scheidegger et al. (2025). The application of spectral deconfounding techniques for Random Forests is novel.
We develop a new algorithm with R-package SDModels (Ulmer & Scheidegger 2025) and show its performance in estimating the direct and unconfounded relationship between the observed covariates and the response. We demonstrate that in presence of hidden confounding, our method, the Spectrally Deconfounded Random Forest (SDForest), outperforms the standard Random Forests in many aspects. If no latent factor or confounding exists, our SDForests and standard Random Forests perform similarly, perhaps with a minimal edge in favor of the classical algorithm. Thus, if one is unsure whether hidden confounding is present, there is much to gain but not much to lose.
1.2 Notation
We denote the largest, the smallest, and $i$-th (non-zero) singular value of any rectangular matrix $A$ by $\lambda_{\max}(A)$, $\lambda_{\min}(A)$ and $\lambda_{i}(A)$ respectively. The condition number is defined as $\textup{cond}(A)\coloneqq\frac{\lambda_{\max}(A)}{\lambda_{\min}(A)}$. Let $\{r_{n}\}_{n=1}^{\infty}$ and $\{k_{n}\}_{n=1}^{\infty}$ be positive constants. We use the notation $k_{n}\coloneqq\Omega(r_{n})$ if $\frac{r_{n}}{k_{n}}=\mathcal{O}(1)$, i.e., if $k_{n}$ has asymptotically at least the same rate as $r_{n}$ and $k_{n}\asymp r_{n}$ if $k_{n}$ and $r_{n}$ have asymptotically the same rate. We write $r_{n}\ll k_{n}$ if $\frac{r_{n}}{k_{n}}=o(1)$.
2 Confounding Model
Throughout this work, we assume the confounding model, written in terms of structural equations
$\begin{split}&X\leftarrow\Gamma^{T}H+E\\ &Y\leftarrow f^{0}(X)+\delta^{T}H+\nu.\end{split}$ (1)
Here, $X\in\mathbb{R}^{p}$ are the predictors, $Y\in\mathbb{R}$ is the response and $H\in\mathbb{R}^{q}$ are unobserved hidden confounding factors. We assume that the confounder influences $X$ with a linear effect $\Gamma\in\mathbb{R}^{q\times p}$ and $Y$ with a linear effect $\delta\in\mathbb{R}^{q}$ (see Appendix C for a note on non-linear confounding); without loss of generality, we can assume $\mathrm{Cov}(H)=I$. Furthermore, $\nu$ is a random variable with mean zero and variance $\sigma_{\nu}^{2}$, and $E$ is a random vector with mean zero and covariance $\Sigma_{E}$, and $E$ and $H$ are uncorrelated. The error term $E$ can be viewed as the unconfounded predictor if $\Gamma$ equals zero. Finally, $f^{0}\in\mathcal{F}$, where $\mathcal{F}$ is some class of functions from $\mathbb{R}^{p}$ to $\mathbb{R}$. The function $f^{0}$ encodes the direct causal relationship of interest, describing the causal relation of $X$ on $Y$. The described model is visualized as a graph in Figure 1.
$X$$Y$$H$$f^{0}$$\Gamma$$\delta$
Figure 1: Confounding model (1), with hidden confounder $H$ affecting $X$ and $Y$ linearly. The function $f^{0}(X)$ encodes the direct effect of $X$ on $Y$.
3 Generic Methodology
We assume that we observe $n$ i.i.d observations from $X$ and $Y$ generated by model (1). We concatenate them row-wise into the design matrix $\mathbf{X}\in\mathbb{R}^{n\times p}$ and the vector of responses $\mathbf{Y}\in\mathbb{R}^{n}$. Classical regression methods ignoring the confounding would estimate $f^{0}$ by minimizing the least squares objective $\hat{f}\coloneqq\text{arg\ min}_{f\in\mathcal{F}}\|\mathbf{Y}-f(\mathbf{X})\|_{2}^{2}$, or a regularized version thereof, for a suitable class $\mathcal{F}$ of functions. This yields an estimate of $\text{arg\ min}_{f\in\mathcal{F}}\mathbb{E}[(Y-f(X))^{2}]=\mathbb{E}[Y|X]=f^{0}(X)+\delta^{T}\mathbb{E}[H|X]$, which is a biased estimate for $f^{0}$ in model (1). We apply a spectral transformation to the least squares objective to remove this confounding bias. Let $Q\in\mathbb{R}^{n\times n}$ be a transformations matrix that depends on the data $\mathbf{X}$. Examples are the trim-transform and the PCA adjustment (Ćevid et al. 2020, Guo et al. 2022, Scheidegger et al. 2025). We use the trim-transform in our empirical results in Section 5, which limits all the singular values of $\mathbf{X}$ to some constant $\tau$, but the methodology can be applied using other spectral transformations. Let $\mathbf{X}=UDV^{T}$ be the singular value decomposition of $\mathbf{X}$, where $U\in\mathbb{R}^{n\times r}$, $D\in\mathbb{R}^{r\times r}$, $V\in\mathbb{R}^{p\times r}$, where $r\coloneqq\min(n,p)$ is the rank of $\mathbf{X}$. The trim transform $Q$ is then defined as
$Q\coloneqq U\tilde{D}U^{T}$ (2)
$\tilde{D}\coloneqq\begin{bmatrix}\tilde{d_{1}}/d_{1}&0&\cdots&0\\ 0&\tilde{d_{2}}/d_{2}&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&\tilde{d_{r}}/d_{r}\end{bmatrix}$
$\tilde{d_{i}}\coloneqq\min(d_{i},\tau)$
with $\tau$ being the median singular value of $\mathbf{X}$, as recommended by Ćevid et al. (2020) (see Appendix B for a visualization). Trimming the first few singular values results in the reduction of the loss in the direction of the first few principal components of $\mathbf{X}$ and in the confounding model (1), this is also the direction containing most of the confounding effects. Thus, reducing this part of the loss results in the reduction of the confounding bias. At the same time, it is very unlikely that the true sparse function $f^{0}(X)$ lies in the direction of the first few principal components unless there is an artificial relation between $f^{0}(.)$ and the covariance matrix of $X$. We will, therefore, minimize a spectrally transformed version of the mean squared error that we refer to as the spectral objective:
$\min_{f\in\mathcal{F}}\frac{\|Q(\mathbf{Y}-f(\mathbf{X}))\|_{2}^{2}}{n}.$ (3)
Theorem 2 below shows that if the confounding follows some assumptions and we have a spectral transformation, we optimize in the limit essentially $\min_{f\in\mathcal{F}}\|Q(f^{0}(\mathbf{X})-f(\mathbf{X})+\nu)\|_{2}^{2}/n$ with $\nu$ as in model (1) and being independent of $X$. This means that we asymptotically minimize a transformed least squares objective without confounding rather than the usual least squares objective with confounding. Figure 2 shows how the spectral transformation changes the correlation between $\mathbf{Y}$ and $f^{0}(\mathbf{X})$ and that we can use $Q\mathbf{Y}$ as an approximation for $Qf^{0}(\mathbf{X})$.
Figure 2: A random realization according to the confounding model (1) with non-linear $f^{0}$ as in (13) and with the same parameter as in Section 5. On the left, we show $f^{0}(\mathbf{X})$ against $\mathbf{Y}$; on the right, the spectrally transformed versions are shown against each other, that is, $Qf^{0}(\mathbf{X})$ versus $Q\mathbf{Y}$. In both visualizations, the line with a slope equal to one, which corresponds to perfect correlation, is shown as a dashed line.
3.1 Assumptions and Technical Motivation
The spectral deconfounding methodology (Ćevid et al. 2020) relies on a set of assumptions that are also crucial for our SDForests. We review these assumptions in the following.
Model:
The data is generated according to the confounding model (1) with
$\mathbb{E}[H]=0\in\mathbb{R}^{q},\;\mathbb{E}[HH^{T}]=I_{q},\;\mathbb{E}[E]=0\in\mathbb{R}^{p},\;\mathbb{E}[HE^{T}]=0\in\mathbb{R}^{p\times q},$ (4)
i.e., $H$ and $E$ are both centered, and they are uncorrelated. The assumption that $H$ has unit covariance matrix is without loss of generality: let $\Sigma_{H}\coloneqq\mathbb{E}[HH^{T}].$ We can then consider the confounding model with $\tilde{H}\coloneqq\Sigma_{H}^{-1/2}H$, $\tilde{\Gamma}\coloneqq\Sigma_{H}^{1/2}\Gamma$ and $\tilde{\delta}\coloneqq\Sigma_{H}^{1/2}\delta$ which satisfies $\mathbb{E}[\tilde{H}\tilde{H}^{T}]=I_{q}$.
Dimensions:
We will see in Theorem 2 below that the confounding effect goes to zero as $\min(n,p)$ grows. Hence, we need to assume that $p$ increases to infinity with $n$. Moreover, we need the number $q$ of confounders to be low-dimensional, i.e., $q\ll\min(n,p)$.
Covariance of $E$:
It is essential that the covariance $\Sigma_{E}\coloneqq\mathbb{E}[EE^{T}]\in\mathbb{R}^{p\times p}$ of the unconfounded part $E$ of $X$ is sufficiently well-behaved. If, for example, $E$ itself had a factor structure, it would be difficult to separate the confounding $\Gamma^{T}H$ from the factor structure in $E$. This well-behavedness assumption is formalized by
$\textup{cond}(\Sigma_{E})=\mathcal{O}(1)\text{ and }\lambda_{\max}(\Sigma_{E})=\mathcal{O}(1).$ (5)
A simple example where (5) is satisfied is when the components of $E$ are uncorrelated and of the same order, i.e., $\Sigma_{E}=\textup{diag}(\sigma_{1}^{2},\ldots,\sigma_{p}^{2})$ with $\max_{i,j=1,\ldots,p}\sigma_{i}^{2}/\sigma_{j}^{2}\leq C_{1}<\infty$ and $\max_{i=1,\ldots,p}\sigma_{i}^{2}\leq C_{2}<\infty$ for some constants $C_{1},C_{2}>0$ independent of $p$. However, more general covariance structures are possible.
Dense confounding:
The dense confounding assumption intuitively means that each component of $H$ affects many components of $X$ and hence is a property of the matrix $\Gamma$. More formally, it is a statement on how large the minimal singular value of $\Gamma$ should be. For Theorem 2 below, we will work under the assumption that
$\lambda_{\min}(\Gamma)=\Omega(\sqrt{p}),$ (6)
although weaker assumptions are possible (Guo et al. 2022, Scheidegger et al. 2025). Equation (6) is, for example, satisfied with high probability if $q/p\to 0$ and either the rows or columns of $\Gamma$ are sampled as i.i.d. sub-Gaussian random vectors with mean zero and covariance $\Sigma_{\Gamma}$ with $\lambda_{\min}(\Sigma_{\Gamma})$ bounded away from zero (see Lemma 6 in Ćevid et al. (2020)).
Spectral transformation:
The spectral transformation $Q$ defined in (2) (trim transform) satisfies
$\lambda_{\max}(Q\mathbf{X})=\mathcal{O}_{P}\left(\sqrt{\max(n,p)}\right).$ (7)
In Guo et al. (2022), (7) is verified for the case where $E$ is a sub-Gaussian random vector and $\lambda_{\max}(\Sigma_{E})=\mathcal{O}(1).$
The following theorem, which is essentially a compilation of results from Ćevid et al. (2020) and Guo et al. (2022), serves as a motivation to construct SDForests based on the spectral objective (3).
Theorem 1.
Assume the confounding model (1) and assume that the conditions (4), (5), (6) and (7) hold. Then, it holds that
$\frac{\|Q(\mathbf{Y}-f(\mathbf{X}))\|_{2}}{\sqrt{n}}=\frac{\|Q(f^{0}(\mathbf{X})-f(\mathbf{X})+\nu)\|_{2}}{\sqrt{n}}+R_{n}$
where $R_{n}=\mathcal{O}_{\mathbb{P}}\left(\frac{\|\delta\|_{2}}{\min(\sqrt{n},\sqrt{p})}\right)$.
In particular, if $\|\delta\|_{2}^{2}\ll\min(n,p)$, we have that $R_{n}=o_{P}(1)$. The condition $\|\delta\|_{2}^{2}\ll\min(n,p)$ holds for example if $q\ll\min(n,p)$ and all the entries of $\delta\in\mathbb{R}^{q}$ are bounded.
4 SDForest Algorithm
In principle, our algorithm works similarly to the original CART algorithm (Breiman et al. 2017) and Random Forests (Breiman 2001). The main difference is that we minimize the spectral objective (3) instead of the classical mean squared error. This results in additional challenges. For example, we can no longer decompose the loss and treat leaves independently but must estimate tree levels globally using linear regression.
4.1 Spectrally Deconfounded Tree
Assume the confounding model (1) with $\mathcal{F}$ being the function class of step functions, e.g.,
$f^{0}(X)\coloneqq\sum_{m=1}^{M}\mathds{1}_{\{X\in R_{m}\}}c_{m},$ (8)
where $(R_{m})_{m=1}^{M}$ are regions dividing the space of $\mathbb{R}^{p}$ into $M$ rectangular parts. Each region has response level $c_{m}\in\mathbb{R}$. We can write the sample version as $f^{0}(\mathbf{X})=\mathcal{P}c$ where $\mathcal{P}\in\{0,1\}^{n\times M}$ is an indicator matrix encoding to which region an observation belongs, i.e. $\mathcal{P}_{i,m}=1$ if the $i$th row of $\mathbf{X}$ belongs to $R_{m}$ and $\mathcal{P}_{i,m}=0$ otherwise. We refer to $\mathcal{P}$ also as a partition, slightly abusing terminology. The vector $c=(c_{1},\ldots,c_{M})\in\mathbb{R}^{M}$ contains the levels corresponding to the different regions. We can estimate $\hat{f}$ in the spirit of (3) with
$(\hat{\mathcal{P}},\hat{c})\coloneqq\text{arg\ min}_{\mathcal{P}^{\prime}\in\{0,1\}^{n\times M},c^{\prime}\in\mathbb{R}^{M}}\frac{\|Q(\mathbf{Y}-\mathcal{P^{\prime}}c^{\prime})\|_{2}^{2}}{n},$ (9)
where $\mathcal{P}^{\prime}$ has to follow a repeated splitting of the space of $\mathbb{R}^{p}$ into rectangular regions. $M$ is fixed here but has to be estimated in practice. $\hat{M}$ also has to be regularized so that it does not go to $n$, and we overfit by having a region for each observation. If the partition $\mathcal{P}$ is known, the spectral objective (3) becomes
$\begin{split}\hat{c}&=\text{arg\ min}_{c^{\prime}\in\mathbb{R}^{M}}\frac{\|Q(\mathbf{Y}-\mathcal{P}c^{\prime})\|_{2}^{2}}{n}\\ &=\text{arg\ min}_{c^{\prime}\in\mathbb{R}^{M}}\frac{\|\mathbf{\tilde{Y}}-\mathcal{\tilde{P}}c^{\prime}\|_{2}^{2}}{n},\end{split}$ (10)
where $\mathbf{\tilde{Y}}\coloneqq Q\mathbf{Y}$ and $\mathcal{\tilde{P}}\coloneqq Q\mathcal{P}$. This is a linear regression problem, and we estimate $c$ by
$\hat{c}=(\mathcal{\tilde{P}}^{T}\mathcal{\tilde{P}})^{-1}\mathcal{\tilde{P}}^{T}\mathbf{\tilde{Y}}.$ (11)
This contrasts the classical CART setting, where, given a partition $\mathcal{P}$, we estimate $c$ with the mean per region $\mu_{m}$. In Equation (10), each $\hat{c}_{m}$ depends not only on the observations belonging to $R_{m}$ but on all the observations due to the rotation and scaling of the spectral transformation. For estimating $\hat{\mathcal{P}}$ we propose Algorithm 1. Algorithm 1 estimates the partition using a tree structure and a repeated splitting of leaves. Since comparing all possibilities for $\mathcal{P}$ is impossible, we grow a tree greedily. Given a current tree, we iterate over all regions and all possible splits. We choose the one that reduces the spectral loss (9) the most, using a subroutine described in Section 4.2, and estimate after each split all the region estimates $\hat{c}$ using Equation (10) in line 17 of Algorithm 1. As the number of leaves grows, the more possible splits we have to compare. In the classical CART setting, an optimal split in a region and its loss decrease remain the same even when another region is divided. Because the spectral transformation $Q$ induces dependency between the observations, the operations for regions depend on all the data. Therefore, to find the next optimal split, we would have to estimate the loss decrease in all potential splits anew in every iteration. Despite this, we reuse the previously estimated loss decrease to save computation time and argue that the change is small and still results in a reasonable split; see also Appendix A (note that we still re-estimate $\hat{c}$ after each iteration). This means that, for each iteration, we only calculate the potential loss decreases for all potential splits in the two new regions. This is repeated until the loss decreases less than a minimum loss decrease after a split. The minimum loss decrease equals a cost-complexity parameter $\mathsf{cp}$ times the initial loss when only an overall mean is estimated. The cost-complexity parameter $\mathsf{cp}$ controls the complexity of a regression tree and acts as a regularization parameter. This should stop $\widehat{M}$ from growing up to $n$. Algorithm 1 results in a regression tree that has been grown aiming to minimize the spectral loss (9) following Equation (10).
4.2 Subroutine
In the $M$th iteration of Algorithm 1, we have the indicator matrix $\hat{\mathcal{P}}=\hat{\mathcal{P}}^{M}\in\{0,1\}^{n\times M}$ with entries $\hat{\mathcal{P}}^{M}_{i,m}=1$ if and only if the $i$th observation lies in region $m$. We encode a candidate split in the region $b$ of the partition using covariate $j$ at the splitting value $s$ by $e\in\{0,1\}^{n}$ where $e$ must be in the support of the $b$th column of $\hat{\mathcal{P}}^{M}$ (i.e. the indices of $1$s in $e$ are a subset of the indices of $1$s of the $b$th column of $\hat{\mathcal{P}}^{M}$)) and the entries depend on $j$ and $s$. A candidate split results then in a candidate partition $\mathcal{P}^{M+1}(e)$ with $M-1$ columns equal to columns $\hat{\mathcal{P}}^{M}$, the $b$th column equal to the $b$th column of $\hat{P}^{M}$ minus $e$ and an additional column equal to $e$. In lines 6-12 of Algorithm 1, we seek to find the optimal $e$ among a large number of candidate splits such that the spectral objective $\|QY-Q\mathcal{P}^{M+1}(e)\hat{c}\|_{2}^{2}$ is minimal, where $\hat{c}$ is the least squares estimator $\hat{c}\coloneqq\arg\min_{c\in\mathbb{R}^{M+1}}\|QY-Q\mathcal{P}^{M+1}(e)c\|_{2}^{2}$. Naively, for every candidate $e$, one would update the indicator matrix $\hat{\mathcal{P}}$, calculate the corresponding least squares estimator $\hat{c}$ and plug it in to obtain the loss decrease. Using the following procedure, the decrease in loss can be calculated more efficiently.
Note that $\textup{span}(Q\mathcal{P}^{M+1}(e))=\textup{span}(Q\hat{\mathcal{P}}^{M},Qe)$. We set $u_{1}^{\prime}\coloneqq Q\cdot(1,\ldots,1)^{T}$ and $u_{1}\coloneqq u_{1}^{\prime}/\|u_{1}^{\prime}\|_{2}$. We proceed by induction and assume that we already have $u_{2},\ldots,u_{M}$ such that $(u_{1},\ldots,u_{M})$ form an orthonormal basis for $\textup{span}(Q\hat{\mathcal{P}}^{M})$. Consequently, $\textup{span}(Q\mathcal{P}^{M+1}(e))=\textup{span}(Q\hat{\mathcal{P}}^{M},Qe)=\textup{span}(u_{1},\ldots,u_{M},Qe)=\textup{span}(u_{1},\ldots,u_{M},u_{M+1}(e))$, where we use orthogonalization to obtain $u_{M+1}(e)$, i.e. $u_{M+1}(e)=u_{M+1}(e)^{\prime}/\|u_{M+1}(e)^{\prime}\|_{2}$ with $u_{M+1}(e)^{\prime}=Qe-\sum_{l=1}^{M}(u_{l}^{T}Qe)u_{l}=(Q-\sum_{l=1}^{M}u_{l}u_{l}^{T}Q)e$.
Let $\Pi_{M+1}(e)\in\mathbb{R}^{n\times n}$ be the orthogonal projection on $\textup{span}(Q\mathcal{P}^{M+1}(e))=\textup{span}(Q\hat{\mathcal{P}}^{M},Qe)=\textup{span}(u_{1},\ldots,u_{M},u_{M+1}(e))$. We seek for $e$ that minimizes $\|Q\mathbf{Y}-Q\mathcal{P}^{M+1}(e)\hat{c}\|_{2}^{2}=\|(I_{n}-\Pi_{M+1}(e))Q\mathbf{Y}\|_{2}^{2}$. But because $(u_{1},\ldots,u_{M},u_{M+1}(e))$ is an orthonormal set, it follows that
$\|(I_{n}-\Pi_{M+1}(e))Q\mathbf{Y}\|_{2}^{2}=\|Q\mathbf{Y}\|_{2}^{2}-\|\Pi_{M+1}(e))Q\mathbf{Y}\|_{2}^{2}=\|Q\mathbf{Y}\|_{2}^{2}-\sum_{l=1}^{M}(u_{l}^{T}Q\mathbf{Y})^{2}-(u_{M+1}(e)^{T}Q\mathbf{Y})^{2}.$
To find the optimal split $e$, it suffices to maximize $\alpha(e)\coloneqq(u_{M+1}(e)^{T}Q\mathbf{Y})^{2}$ among the candidate splits. Once the optimal split $e^{\ast}$ is found, one can define $u_{M+1}\coloneqq u_{M+1}(e^{\ast})$ and $\hat{\mathcal{P}}^{M+1}\coloneqq\mathcal{P}^{M+1}(e^{\ast})$ and proceed with step $M+2$.
Algorithm 1 Spectrally Deconfounded Regression Tree
1:Inputs:
2:      $\mathbf{X}\in\mathbb{R}^{n\times p}$, $\mathbf{Y}\in\mathbb{R}^{n}$, $Q\in\mathbb{R}^{n\times n}$, $\mathsf{cp}\in\mathbb{R}$, $M_{max}\in\mathbb{N}$
3:Initialize:
4:      $M\leftarrow 1$ $\hat{\mathcal{P}}\leftarrow(1,\ldots,1)^{T}\in\mathbb{R}^{n\times 1}$ $\tilde{\hat{\mathcal{P}}}\leftarrow Q\hat{\mathcal{P}}$ $\mathbf{\tilde{Y}}\leftarrow Q\mathbf{Y}$ $u\leftarrow\hat{\mathcal{P}}/\|\hat{\mathcal{P}}\|_{2}$ $Q^{d}\leftarrow Q-uu^{T}Q$ $\hat{c}\in\mathbb{R}^{M}\leftarrow\text{arg\ min}_{c^{\prime}\in\mathbb{R}^{m}}\|\mathbf{\tilde{Y}}-\tilde{\hat{\mathcal{P}}}c^{\prime}\|_{2}^{2}/n$ $l^{init}\leftarrow l^{temp}\leftarrow\|\mathbf{\tilde{Y}}-\tilde{\hat{\mathcal{P}}}\hat{c}\|_{2}^{2}/n$ $l^{dec}\in\mathbb{R}^{M}\leftarrow 0$ $\mathcal{B}\leftarrow 1$
5:for $M=1$ to $M_{max}$ do
6:     for $b$ in $\mathcal{B}$ do $\triangleright$ subroutine
7:         for $(j,s)$ in potential splits in region $b$ do
8:              $e_{b,j,s}\in\{0,1\}^{n}\leftarrow$ indices of samples belonging to the new partition
9:              $u\leftarrow Q^{d}e_{b,j,s}/\|Q^{d}e_{b,j,s}\|_{2}$
10:              $\alpha_{b,j,s}\leftarrow(u^{T}\mathbf{\tilde{Y}})^{2}$
11:         end for
12:     end for
13:     $(b^{*},j^{*},s^{*})\leftarrow\text{arg\ max }{\alpha_{b,j,s}}$triangleright optimal split over $b\in\{1,\ldots,M\}$
14:     $u\leftarrow Q^{d}e_{b^{*},j^{*},s^{*}}/\|Q^{d}e_{b^{*},j^{*},s^{*}}\|_{2}$
15:     $Q^{d}\leftarrow Q^{d}-uu^{T}Q$
16:     $\hat{\mathcal{P}}^{*}\leftarrow$ splitting $\hat{\mathcal{P}}$ at $(b^{*},j^{*},s^{*})$triangleright resulting in $\hat{\mathcal{P}}^{*}\in\mathbb{R}^{n\times(M+1)}$
17:     $\hat{c}^{*}\leftarrow\text{arg\ min}_{c^{\prime}\in\mathbb{R}^{M}}\|\mathbf{\tilde{Y}}-\tilde{\hat{\mathcal{P}}}^{*}c^{\prime}\|_{2}^{2}/n$
18:     $l^{*}\leftarrow\|\mathbf{\tilde{Y}}-\tilde{\hat{\mathcal{P}}}^{*}\hat{c}^{*}\|_{2}^{2}/n$
19:     $d\leftarrow l^{temp}-l^{*}$
20:     if $d>\mathsf{cp}*l^{init}$ then
21:         $\hat{\mathcal{P}}\leftarrow\hat{\mathcal{P}}^{*}$
22:         $\hat{c}\leftarrow\hat{c}$
23:         $l^{temp}\leftarrow l^{*}$
24:         $\mathcal{B}\leftarrow(b^{*},M+1)$triangleright the new partitions
25:     else
26:         break
27:     end if
28:end for
4.3 Spectrally Deconfounded Random Forests
The next natural step is to utilize spectrally deconfounded regression trees (SDTree) to construct spectrally deconfounded Random Forests (SDForests) for estimating arbitrary functions. Random Forests were introduced by Breiman (2001) and have been successfully employed in numerous applications. The idea is to combine multiple regression trees into an ensemble to decrease variance and obtain a smoother function. Ensembles work best if the different models are independent of each other. To decorrelate the regression trees as much as possible from each other, we have two mechanisms. The first one is bagging (Breiman 1996), where we train each regression tree on an independent bootstrap sample of the observations, i.e., we draw a random sample of size $n$ with replacement from the observations. The second mechanic to decrease the correlation is that only a random subset of the covariates is available for each split. Before each split, we sample $\mathsf{mtry}\leq p$ from all the covariates and choose the one that reduces the loss the most from those. Since we already have Algorithm 1, it only takes minor changes to build an SDForest with it. In Algorithm 1 before line 13, we sample a set $p_{\mathsf{mtry}}$ of size $\mathsf{mtry}$ from the covariates and in line 13, we only check in this set of randomly sampled covariates for the best split, e.g., $j\in p_{\mathsf{mtry}}$. This procedure gives the spectrally deconfounded Random Forest in Algorithm 2. We predict with all the trees separately and use the mean over all trees as the Random Forest prediction. This gives the estimated function
$\hat{f}(X)\coloneqq\frac{1}{N_{tree}}\sum_{t=1}^{N_{tree}}SDTree_{t}(X).$ (12)
Algorithm 2 Spectrally Deconfounded Random Forest
Inputs:
      $\mathbf{X}\in\mathbb{R}^{n\times p}$, $\mathbf{Y}\in\mathbb{R}^{n}$, $Q\in\mathbb{R}^{n\times n}$, $N_{tree}\in\mathbb{N}$, $\mathsf{mtry}\in[1,p]$
for $t=1$ to $N_{tree}$ do
     $X^{t}\leftarrow$ bootstrap sample of $X$
     $SDtree_{t}\leftarrow$ SDTree from Algorithm 1 with random set of covariates of size $\mathsf{mtry}$ at each split using $X^{t}$
end for
5 Empirical Results
For the simulation study, we simulate data according to the confounding model (1) with a random $f^{0}$ using the Fourier basis
$f^{0}(X)\coloneqq\sum_{j=1}^{p}\mathds{1}_{\{j\in\mathcal{J}_{s}\}}\sum_{k=1}^{K}a_{j,k}\cos(0.2k\cdot x_{j})+b_{j,k}\sin(0.2k\cdot x_{j})$ (13)
where $\mathcal{J}_{s}$ is a random subset of ${1,\ldots,p}$ of size four being the parents of $Y$ (among the $X$ variables). We simulate with $n=1000$, $p=500$, and $q=20$. The entries of $\mathbf{E}\in\mathbb{R}^{n\times p}$, $\mathbf{H}\in\mathbb{R}^{n\times q}$, $\delta\in\mathbb{R}^{q}$, and $\Gamma\in\mathbb{R}^{q\times p}$ are sampled i.i.d. from a Gaussian with expectation zero and $\sigma=1$. The additional noise $\mathbf{\nu}\in\mathbb{R}^{n}$ is sampled i.i.d. from a Gaussian with mean zero and $\sigma_{\nu}=0.1$. For the four causal parents, we sample the coefficients $a_{j,k}$ and $b_{j,k}$ uniformly on $[-1,1]$ with the number of basis functions fixed at $K=2$ for the additive function.
We use the SDForest with a hundred trees, $\mathsf{mtry}=\lfloor 0.5p\rfloor$ and $Q$ as the trim-transform (2) (Ćevid et al. 2020) to estimate the causal function. We compare the results of the SDForest to the estimated function by the classical Random Forest using the r package ranger (Wright & Ziegler 2017).
In Figure 3, we compare the variable importance of the SDForest and the classical Random Forest for one simulation run. The variable importance of a covariate is calculated as the sum of the loss decrease resulting from all splits that divide a region using that covariate. The mean of the variable importance of all trees results in the variable importance for the forest. For the SDForest, we report the reduction of the spectral loss (3) instead of the MSE. For the SDForest, the four most important covariates are also the four true causal parents of $Y$. Three of those also have a clear separation from the remaining 497 covariates. For the classical Random Forest, the variable importance for the true causal parents lies within the bulk of spurious covariates. The most important covariate among the true causal parents is only on rank 102 and we would have to use almost all the covariates to include all the causal parents in the selected set.
Figure 3: Comparison of variable importance for a realization of model (13) between the classical Random Forest estimated by ranger and the SDForest. The variable importance for both methods is scaled to the interval $[0,1]$ and log-transformed. The true causal parents of the response $Y$ are marked as crosses.
Instead of examining the variable importance of the fully grown trees in the SDForest, we can also examine the regularization paths of the covariates. Figure 4 shows on the left side the variable importance for the SDForest against increasing regularization, where one increases $\mathsf{cp}$ subsequently pruning the trees. Here again, three of the causal parents appear. On the right side in Figure 4, we show stability selection (Meinshausen & Bühlmann 2010), where $\Pi$ is the ratio of trees that use a particular covariate in the forest given increasing regularization. In the stability selection paths, we also see the fourth causal parent.
Figure 4: Regularization paths of the SDForest estimated on a realization of model (13) when varying the cost-complexity parameter $\mathsf{cp}$ resulting in more or less pruned trees. Each curve corresponds to a single covariate. On the left side are the variable importance paths for different strengths of regularization shown. On the right side are the stability selection paths against the strength of regularization shown. $\Pi$ corresponds to the ratio of trees in the forest that use a covariate. The truly causal parents of the response $Y$ correspond to the darker, thicker lines.
In addition to screening for the sparse causal set among a large number of covariates, we can also look at the functional dependence of the response $Y$ on the causal parents. We use partial dependence plots (Friedman 2001) to visualize the partial dependence. The idea is to predict $\hat{f}(\mathbf{X})$ for each observation and vary $\mathbf{X}_{j}$ on an interval while we keep the other covariates at the observed values. This gives for every observation a different function of $X_{j}$. The mean over all the observations can then be shown as a representative marginal effect of $X_{j}$ on the response $Y$. Figure 5 shows how the SDForest estimates the true causal function. Especially for covariate 34, the estimated function approximates the true function well. Covariate 108, the covariate with only slightly higher importance than the bulk of spurious covariates, has almost a constant influence on $Y$. This shows some limitations of estimating a sparse causal relationship in the presence of hidden confounding and high dimensionality. Such a nuanced true function as for variable 108 might be hard to estimate with the additional disturbance of confounding and noise. The classical Random Forest distributes the estimated function among all the spurious covariates and estimates almost a constant function for all four causal parents. The partial dependence plots for the estimated classical Random Forest are shown in Figure 6.
Figure 5: Partial dependence plots of the estimated SDForest for the four true causal parents of the response $Y$. The dashed line is the corresponding true partial causal function. The light lines show the observed empirical partial functions for 19 randomly selected observations, and the thick line is the average of all observed partial functions.
Figure 6: Partial dependence plots as in Figure 5 but now with the estimated classical Random Forest by ranger for the four true causal parents of the response $Y$. The classical Random Forest essentially leads to constant functions, whereas the true dashed lines are varying.
5.1 Performance of SDForests
To quantify the performance of the SDForests in estimating the causal function, we perform simulation studies varying the different dimensions in the simulation. The default dimensions are $n=500$, $p=500$, and $q=20$, and we randomly draw data according to model (13). Each of those dimensions is varied separately to estimate the dependence of the performance on these different factors. Every experiment is repeated 200 times, where every time, the whole data-generating process is newly drawn at random. The performance is measured by the mean distance of the true causal function and the estimated function evaluated at 500 test observations $f_{mse}\coloneqq\frac{1}{n_{test}}\sum_{i=1}^{n_{test}}(f^{0}(x_{test,i})-\hat{f}(x_{test,i}))^{2}$. The performance of the SDForests and the classical Random Forests depending on different dimensions is shown in Figure 7. Theorem 2 shows how the spectral objective (3) behaves as the dimensionality $p$, and the number of observations $n$ grows to infinity. The term $R_{n}$ goes asymptotically with $\min(\sqrt{n},\sqrt{p})$ to zero. So, we need both $n$ and $p$ to grow for consistency. With this simulation study, we look at the dependence of the performance on those quantities in practice. Figure 7a shows the error distribution depending on the sample size. The SDForests perform clearly better in estimating the causal function, and the error decreases with a larger sample size. We see the dependence of the performance on the number of covariates in Figure 7b. At $p>20$, we see how the SDForests start to perform significantly better than the classical Random Forests. When increasing $p$ above 50, the error no longer seems to decrease, which hints that there is a certain threshold of $p$ after which deconfounding is successful. Even with only the four causal parents as covariates, the SDForests do not perform worse than the classical versions, and we really do not seem to lose anything with SDForests, even in low-dimensional settings. With stronger confounding, estimating the causal function becomes increasingly difficult. Not only does the bias increase but so does the variance in the response. Therefore, we expect the error to increase with increasing confounding even for the SDForests. Figure 7c shows this behavior with a simulation study, where we increase the number of confounding variables. For $q=0$, the setting corresponds to the classical model without confounding, and both the classical Random Forests and the SDForests perform really well. It is important to note that we do not lose much, even when the data is not confounded, by applying the spectral deconfounding. However, as can be seen for increasing confounding, we can gain a lot if there is hidden confounding present. In Figure 7d, we follow up on the assumption of dense confounding. Here, instead of having random effects of the confounders on all the covariates, we simulate the data with only a random subsample of the covariates being affected by confounding. The denser the confounding becomes, the better SDForests perform. With around 200 covariates affected (40%), the SDForests start performing similarly as if there were no confounding, while with sparser confounding, SDForest still outperforms ranger.
Figure 7: Mean squared error of the estimated causal function $\hat{f}(X)$ by classical Random Forests estimated by ranger and the SDForests depending on different simulation parameters. In subfigure a), we show how the performance depends on the sample size. In subfigure b), we show how the performance depends on the dimensionality of the observed data. The dependence of the performance on the amount of confounding is shown in subfigure c), where zero confounders corresponds to the classical setting without confounding. Subfigure d) shows the dependency of the performance on the number of affected covariates by the confounding, investigating the importance of dense confounding. Both algorithms estimate a hundred trees using $\mathsf{mtry}=\lfloor 0.5p\rfloor$.
6 Single-Cell Data
We apply SDForest and ranger to Single-cell gene expression data to compare the resulting model on the scRNA-seq dataset for the cell RPE1 generated by Replogle et al. (2022). We use the preprocessed and filtered dataset provided by Chevalley et al. (2025). As the response, we choose the gene for EIF1 following the arguments of Shen et al. (2023) conjecturing that EIF1 might be a leaf node in the causal graph of genes. Using the observational data without any interventional gene knockouts ($n=11485$ observations), we use all the other gene expressions ($p=382$) as predictors and fit both ranger and SDForest. For both methods, we fit 500 trees using default parameters. To test for robustness against dense confounding, we synthetically perturb the original cleaned and filtered dataset. For this, we construct $\mathbf{X}_{\tau}\coloneqq\mathbf{X}+\mathbf{H}\Gamma\tau$ and $\mathbf{Y}_{\tau}\coloneqq Y+\mathbf{H}\delta\tau$, where $\mathbf{Y}$ is the original expression of EIF1 and $\mathbf{X}$ are the other original gene expressions used for prediction. The entries of $\mathbf{H}\in\mathbb{R}^{n}$, $\Gamma\in\mathbb{R}^{1\times 382}$, and $\delta\in\mathbb{R}$ are sampled i.i.d. from $\mathcal{N}(0,1)$ while we vary $\tau$ to increase the added dense confounding. We fit both methods to $\mathbf{X}_{\tau}$ and $\mathbf{Y}_{\tau}$ with increasing $\tau$ and analyze how much the estimated functions change. The change in the estimated functions is calculated as the change of the out-of-bag predictions of $f^{\text{\emph{ranger}}}_{\tau}(\mathbf{X}^{\tau})$ and $f^{\text{SDF}}_{\tau}(\mathbf{X}^{\tau})$ depending on $\tau$
$\frac{\|f^{\text{method}}_{\tau}(\mathbf{X}_{\tau})-f^{\text{method}}_{0}(\mathbf{X}_{0})\|_{2}^{2}}{n}.$
We repeat the synthetic confounding of the data by sampling 20 times $\mathbf{H}$, $\Gamma$, and $\delta$ and increasing $\tau$. The distributions of the function changes are shown in Figure 8. With no added confounding, the difference between the two functions is small (see caption of Figure 8). This hints at a finding that there is no major dense confounding present. As we increasingly add dense confounding, we clearly see how the estimated function by ranger changes, while SDForest demonstrates robustness against this added perturbation.
Figure 8: Change in prediction $\frac{\|f^{\text{method}}_{\tau}(\mathbf{X}_{\tau})-f^{\text{method}}_{0}(\mathbf{X}_{0})\|_{2}^{2}}{n}$ of ranger and SDForest when perturbing the data with additional dense confounding. At $\tau$ = 0, we show the initial difference in prediction of ranger and SDForest when estimating and predicting on the unperturbed data $\frac{\|f^{\text{SDF}}_{0}(\mathbf{X}_{0})-f^{\text{\emph{ranger}}}_{0}(\mathbf{X}_{0})\|_{2}^{2}}{n}=0.0016$.
7 Conclusion
We propose the Spectrally Deconfounded Random Forest algorithm SDForest with R-package SDModels (Ulmer & Scheidegger 2025) to estimate direct regression functions in high-dimensional data in the presence of dense hidden confounding. This can be used to screen for relevant covariates among a large set of variables, and the procedure provides robustness against unobserved confounding, gaining much in the confounded case but losing little if no confounding is present. SDModels provides functions such as regularization paths and stability selection to screen for relevant covariates as well as partial dependence plots to understand their relation to the variable of interest. We demonstrate the empirical behavior of SDForests in various settings, including challenging cases with sparse confounding, to illustrate some fundamental limitations (which are not found to be clearly worse than those of classical Random Forests).
Many potentially confounded high-dimensional applications are about classification instead of regression. Currently, SDForest is only applicable for regression tasks, and it is important to extend this methodology to classification. Another open question is whether one can combine spectral deconfounding with Quantile Regression Forests (Meinshausen 2006) or Distributional Random Forests (Hothorn & Zeileis 2021, Ćevid et al. 2022) to gain access to prediction intervals or constructing confidence intervals using techniques as in Guo et al. (2022) for the linear case and Näf et al. (2023) using Random Forests.
Supplementary Materials
Appendix A: Approximation of splitting criteria
Appendix B: Visualization of Spectral Transformation of singular values
Appendix C: Notes on Non-linear Confounding
Appendix D: Proofs of all theoretical results
Code: All the code used for this paper is available here: https://github.com/markusul/SDForest-Paper
R-package for SDForests: R-package SDModels provides software for non-linear spectrally deconfounded models: https://github.com/markusul/SDModels
Acknowledgment
We thank Mathieu Chevalley for his support in applying our method to the single-cell dataset.
Funding
M. Ulmer and C. Scheidegger gratefully acknowledge financial support from the Swiss National Science Foundation, grant no. 214865.
Disclosure statement
The authors report there are no competing interests to declare
SUPPLEMENTARY MATERIAL
Appendix A Approximation of splitting criteria
In the $M$th iteration of Algorithm 1, we have to find in each region the split that reduces the loss the most. This means we have $M$ regions in each iteration, each with an optimal split to choose from. When we split the region with the optimal split, we have two new regions. Now, due to the spectral transformation, the samples, as well as the different regions containing a subset of the samples, are not independent. Therefore, to truly find the next best split, we would need to determine the optimal split and its corresponding loss decrease in each region anew. This is done in Figure 9 using the method SDT2. In our studies and as the default in the R-package, we only estimate the optimal split and its loss decrease for the two new regions and reuse all the other estimates from the previous iteration. This saves substantial computational time and we argue that a previously good split stays reasonable. We compare the performance of the two options in Figure 9. We simulate data following the confounding model (1) and with the same parameter as in Section 5 but using a random regression tree for $f^{0}$ as in Equation 8. The random regression tree is grown using random splits until there are ten leaf nodes. While the reestimation of all splits in SDT2 might be a bit better, we do not see a significant decrease in performance when using the computationally much more efficient approximation SDT1. In the case of smooth underlying regression functions, we expect to see similar behavior (because we can consider the approximation error of a smooth function with a tree that remains unaffected and the estimation error of the best-approximated tree function, which is analogous to the discussion above).
Figure 9: Mean squared error of the estimated causal function $\hat{f}(X)$ by SDTree using $\mathsf{cp}=0.01$. SDT1 corresponds to Algorithm 1 while SDT2 uses $\mathcal{B}\leftarrow(1,\ldots,M+1)$ in line 22 instead of only the new partitions. The data is simulated using the confounding model (1) and with the same parameter as in Section 5 but using a random regression tree for $f^{0}$ as in Equation 8.
Appendix B Transformation of singular values
The spectral transformation described in Section 3 shrinks the first few singular values to decrease confounding bias. This is visualized in Figure 10 for the PCA adjustment and the trim-transform. PCA adjustment removes all the signal from the first 20 principal components (we assume here that the number of 20 hidden factors is known). This should not reduce the signal of $f^{0}(X)$ as it lies in the span of a sparse set of covariates. However, choosing the right number of principal components in real data is subtle (Owen & Wang 2016) and may result in unwanted removal of signal of $f^{0}(X)$. The trim-transform is much more insensitive to the problem with the unknown number of confounders, as it only limits the top half of the singular values to their median and does not completely remove any principal components and their associated signal completely.
Figure 10: Singular values of $\mathbf{X}$, $Q_{trim}\mathbf{X}$, and $Q_{pca}\mathbf{X}$. The data is simulated using the confounding model (1) with the same parameters as in Section 5 ($q=20$ hidden factors), but with $p=100$ to increase visibility.
Appendix C Non-linear Confounding
In Equation 1 and Section 3.1, we assumed that the hidden confounder affects both the $X$ and $Y$ linearly. This assumption is not testable and might not hold in practice. We present here some heuristic arguments for when and why spectral deconfounding could still work well with nonlinear confounding. Without loss of generality, assume that $H$ is univariate and assume that the data is generated according to
$Y=f(X)+d(H)+\nu,\quad X_{j}=g_{j}(H)+E_{j},\,j=1,\ldots,p,$ (14)
for some (potentially) nonlinear functions $d$ and $g_{j}$, $j=1,\ldots,p$ from $\mathbb{R}\to\mathbb{R}$. Assume that $g_{j}(\cdot),j=1,\ldots,p$ and $d(\cdot)$ can be well-approximated by a common set of basis functions, $(b_{k}(\cdot))_{k=1}^{K}$, i.e., $g_{j}(\cdot)\approx\sum_{k=1}^{K}\Gamma_{k,j}b_{k}(\cdot)$, $j=1,\ldots,p$ and $d(\cdot)\approx\sum_{k=1}^{K}\delta_{k}b_{k}(\cdot)$. Let $B=(b_{1}(H),\ldots,b_{K}(H))^{T}\in\mathbb{R}^{K}$. Then, it approximately holds that
$Y\approx f(X)+\delta^{T}B+\nu,\quad X\approx\Gamma^{T}B+E,$
i.e., we are approximately in the setting of Equation 1 with linear confounding with $H$ being replaced by $B$ (and without loss of generality, by orthogonalization, we can assume $\mathrm{Cov}(B)=I$). If $\Gamma$ is dense, it is reasonable that spectral deconfounding still works well. Intuitively, $\Gamma$ will be dense if the $g_{j}$ are all “sufficiently different”. Moreover, $d(\cdot)$ should be “similar” to the $g_{j}$ such that there is an approximation with a common basis. The number $K$ in the basis approximations is then analogous to the number of confounding variables. If we assume that the functions $d(\cdot)$ and $g_{j}(\cdot)$, $j=1,\ldots,p$ are not too complicated, it is still reasonable to assume that $K$ is small and we do not need to know it.
Empirically, we simulate data similar to Section 5. In addition to the non-linear function $f^{0}(X)$, we simulate non-linear confounding using Equation 14 where we simulate $g_{j}$ and $d$ using different random functions using the Fourier basis (also using Equation 13). For the confounding effect on $X$ and $Y$, we use $K=12$ basis functions and sample all the coefficients uniformly on $[-1,1]$ for the effect on $X$ and on $[-2,2]$ for the effect on $Y$. Here, we use $q=1$, $p=300$, $n=500$, and let only one covariate affect the response $Y$. For reasonable noise level and confounding strength, we sample $\delta\in\mathbb{R}$ i.i.d. from a Gaussian with mean zero and $\sigma=2$ and the additional noise $\nu\in\mathbb{R}^{n}$ from a Gaussian with mean zero and $\sigma_{\nu}=0.01$. All the other parameters in the simulation stay the same as in Section 5. In Figure 11, we show the singular values of $\mathbf{X}$ of a random realization. We observe that six instead of just one singular value spike due to the non-linear confounding. Apparently, this number is lower than the number of basis functions $K=12$, and the spiking effect is only visible for the first six components. In this setting, $\mathbf{Y}$ is an even worse approximation for $f^{0}(\mathbf{X})$, but the spectral transformation still results in a clear correlation between $Q\mathbf{Y}$ and $Qf^{0}(\mathbf{X})$, see Figure 12. In Figure 13, we show the dependence of $\mathbf{Y}$, $f^{0}(\mathbf{X})$, and $f^{SDForest}(\mathbf{X})$ on the single causal parent $X_{80}$. The observations (points in the figure) show a complicated dependency to $X_{80}$. At the same time, the estimated relationship by SDForest, represented by the thick line, stays close to the true causal function (dashed line). We repeat this simulation a hundred times and report the test performance, using 500 data points, of classical Random Forests and the SDForest in estimating the true function $f^{0}(\mathbf{X})$. The distribution of these performances is shown in Figure 14, where we clearly see that SDForest outperforms the classical Random Forests in this specific non-linear confounding setting as well.
Figure 11: Singular values of a random realization of $\mathbf{X}$ affected non-linearly by a single confounder.
Figure 12: A random realization of the non-linearly confounded process. On the left, we show $f^{0}(\mathbf{X})$ against $\mathbf{Y}$; on the right, the spectrally transformed versions are shown against each other, that is, $Qf^{0}(\mathbf{X})$ versus $Q\mathbf{Y}$. In both visualizations, the line with a slope equal to one, which corresponds to perfect correlation, is shown as a dashed line. This is the same visualization as in Figure 2 for the linear confounding.
Figure 13: Partial dependence plots of the estimated SDForest for the true causal parent of the response $Y$ in the non-linearly confounded setting. The dashed line is the corresponding true partial causal function. The light lines show the observed empirical partial functions for 19 randomly selected observations, and the thick line is the average of all observed partial functions.
Figure 14: Mean squared error of the estimated causal function $\hat{f}(X)$ by classical Random Forests estimated by ranger and the SDForests in the non-linearly confounded setting.
Appendix D Proofs
Theorem 2.
Assume the confounding model (1) and assume that the conditions (4), (5), (6) and (7) hold. Then, it holds that
$\frac{\|Q(\mathbf{Y}-f(\mathbf{X}))\|_{2}}{\sqrt{n}}=\frac{\|Q(f^{0}(\mathbf{X})-f(\mathbf{X})+\nu)\|_{2}}{\sqrt{n}}+R_{n}$
where $R_{n}=\mathcal{O}_{\mathbb{P}}\left(\frac{\|\delta\|_{2}}{\min(\sqrt{n},\sqrt{p})}\right)$.
Proof.
As in Guo et al. (2022) and Ćevid et al. (2020), let $b=\arg\min_{b^{\prime}\in\mathbb{R}^{p}}\mathbb{E}[(X^{T}b^{\prime}-H^{T}\delta)^{2}]=\mathbb{E}[XX^{T}]^{-1}\mathbb{E}[XH^{T}\delta]$, i.e. $X^{T}b$ is the $L_{2}$ projection of $H^{T}\delta$ onto $X$. By (1) and the triangle inequality, we have that
$\begin{split}&\left|\frac{\|Q(\mathbf{Y}-f(\mathbf{X}))\|_{2}}{\sqrt{n}}-\frac{\|Q(f^{0}(\mathbf{X})-f(\mathbf{X})+\nu)\|_{2}}{\sqrt{n}}\right|\\ &\leq\frac{\|Q(\mathbf{Y}-f(\mathbf{X}))-(Q(f^{0}(\mathbf{X})-f(\mathbf{X})+\nu))\|_{2}}{\sqrt{n}}\\ &=\frac{\|Q(f^{0}(\mathbf{X})+\mathbf{H}\delta+\nu-f(\mathbf{X}))-(Q(f^{0}(\mathbf{X})-f(\mathbf{X})+\nu))\|_{2}}{\sqrt{n}}\\ &=\frac{\|Q(f^{0}(\mathbf{X})-f(\mathbf{X})+\mathbf{X}b+(\mathbf{H}\delta-\mathbf{X}b)+\nu)-(Q(f^{0}(\mathbf{X})-f(\mathbf{X})+\nu))\|_{2}}{\sqrt{n}}\\ &\leq\frac{\|Q\mathbf{X}b\|_{2}}{\sqrt{n}}+\frac{\|Q(\mathbf{H}\delta-\mathbf{X}b)\|_{2}}{\sqrt{n}}\end{split}$
From Lemma 1 below, we have that $\|b\|_{2}=\mathcal{O}(\|\delta\|_{2}/\sqrt{p})$ and hence using (7)
$\frac{\|Q\mathbf{X}b\|_{2}}{\sqrt{n}}\leq\frac{1}{\sqrt{n}}\lambda_{\max}(Q\mathbf{X})\|b\|_{2}=\mathcal{O}\left(\|\delta\|_{2}\frac{\max(\sqrt{n},\sqrt{p})}{\sqrt{np}}\right)=\mathcal{O}\left(\frac{\|\delta\|_{2}}{\min(\sqrt{n},\sqrt{p})}\right).$
By Lemma 2 below, the second term behaves as
$\frac{1}{\sqrt{n}}\|Q(\mathbf{H}\delta-\mathbf{X}b)\|_{2}=\mathcal{O}_{\mathbb{P}}\left(\frac{\|\delta\|_{2}}{\sqrt{p}}\right),$
which concludes the proof. ∎
Lemma 1 (Parts of Lemma 6 in Ćevid et al. (2020)).
Assume that the confounding model (1) satisfies the assumptions (4), (5) and (6). Then we have,
$\|b\|_{2}^{2}=\|\mathbb{E}[XX^{T}]^{-1}\mathbb{E}[XH^{T}]\delta\|_{2}^{2}\leq\textup{cond}(\Sigma_{E})\cdot\frac{\|\delta\|_{2}^{2}}{\lambda_{\min}(\Gamma)^{2}}=\mathcal{O}\left(\frac{\|\delta\|_{2}^{2}}{p}\right).$
Lemma 2.
Under the conditions of Theorem 2,
$\frac{1}{n}\|Q(\mathbf{H}\delta-\mathbf{X}b)\|_{2}^{2}=\mathcal{O}_{\mathbb{P}}\left(\frac{1}{p}\right)$
Proof.
Observe that $\frac{1}{n}\|Q(\mathbf{H}\delta-\mathbf{X}b)\|_{2}^{2}\leq\frac{1}{n}\|Q\|_{op}^{2}\|\mathbf{H}\delta-\mathbf{X}b\|_{2}^{2}$, so it suffices to show that $\frac{1}{n}\|\mathbf{H}\delta-\mathbf{X}b\|_{2}^{2}=\mathcal{O}_{\mathbb{P}}\left(\frac{1}{p}\right)$. By Markov’s inequality, it suffices to show
$\mathbb{E}\left[\frac{1}{n}\|\mathbf{H}\delta-\mathbf{X}b\|_{2}^{2}\right]=\mathcal{O}\left(\frac{1}{p}\right).$ (15)
For this, we follow the arguments of Guo et al. (2022). Our term $\mathbb{E}\left[\frac{1}{n}\|\mathbf{H}\delta-\mathbf{X}b\|_{2}^{2}\right]$ corresponds to $\Delta$ in (47) in A.4. there. We can follow the proof of (35) in Lemma 2 given in Section C.4 in Guo et al. (2022). For this, we write
$\frac{1}{n}\|\mathbf{H}\delta-\mathbf{X}b\|_{2}^{2}=\frac{1}{n}\sum_{i=1}^{n}(\mathbf{H}_{i}^{T}\delta-\mathbf{X}_{i}^{T}b)^{2}$
Hence,
$\begin{split}\mathbb{E}\left[\frac{1}{n}\|\mathbf{H}\delta-\mathbf{X}b\|_{2}^{2}\right]&=\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}[(\mathbf{H}_{i}^{T}\delta-\mathbf{X}_{i}^{T}b)^{2}]\\ &=\mathbb{E}[(H^{T}\delta-X^{T}b)^{2}]\\ &=\mathbb{E}[\delta^{T}HH^{T}\delta-2\delta^{T}HX^{T}b+b^{T}XX^{T}b]\\ &=\delta^{T}\mathbb{E}[HH^{T}]\delta-2\delta^{T}\mathbb{E}[HX^{T}]b+b^{T}\mathbb{E}[XX^{T}]b\\ &=\delta^{T}\mathbb{E}[HH^{T}]\delta-2\delta^{T}\mathbb{E}[HX^{T}]\mathbb{E}[XX^{T}]^{-1}\mathbb{E}[XH^{T}]\delta+\delta^{T}\mathbb{E}[HX^{T}]\mathbb{E}[XX^{T}]^{-1}\mathbb{E}[XH^{T}]\delta\\ &=\delta^{T}\mathbb{E}[HH^{T}]\delta-\delta^{T}\mathbb{E}[HX^{T}]\mathbb{E}[XX^{T}]^{-1}\mathbb{E}[XH^{T}]\delta\\ &=\delta^{T}(I_{q}-\Gamma(\Gamma^{T}\Gamma+\Sigma_{E})^{-1}\Gamma^{T})\delta\end{split}$
where we used the definition of $b$ and (4). As in equation (134) in the supplementary materials in Guo et al. (2022), using Woodbury’s identity, we have
$I_{q}-\Gamma(\Gamma^{T}\Gamma+\Sigma_{E})^{-1}\Gamma^{T}=(I_{q}+\Gamma\Sigma_{E}^{-1}\Gamma^{T})^{-1}.$
Let $C=\Sigma_{E}^{-1/2}\Gamma^{T}$ and $C=U_{C}D_{C}V_{C}^{T}$ be the singular values decomposition of $C$. Then, we have
$\begin{split}\mathbb{E}\left[\frac{1}{n}\|\mathbf{H}\delta-\mathbf{X}b\|_{2}^{2}\right]&=\delta^{T}(I_{q}+\Gamma\Sigma_{E}^{-1}\Gamma^{T})^{-1}\delta\\ &\leq\|\delta\|_{2}^{2}\lambda_{\max}\left((I_{q}+\Gamma^{T}\Sigma_{E}^{-1}\Gamma)^{-1}\right)\\ &=\|\delta\|_{2}^{2}\lambda_{\max}\left((I_{q}+C^{T}C)^{-1}\right)\\ &=\|\delta\|_{2}^{2}\lambda_{\max}\left(V_{C}(I_{q}+D_{C}^{T}D_{C})^{-1}V_{C}^{T}\right)\\ &=\|\delta\|_{2}^{2}(1+\lambda_{\min}(D_{C}^{T}D_{C}))^{-1}\\ &=\|\delta\|_{2}^{2}(1+\lambda_{\min}(C^{T}C))^{-1}\\ &\leq\|\delta\|_{2}^{2}\lambda_{\min}(C^{T}C)^{-1}\end{split}$
Note that
$\begin{split}\lambda_{\min}(C^{T}C)&=\min_{x\neq 0}\frac{x^{T}C^{T}Cx}{x^{T}x}\\ &=\min_{x\neq 0}\frac{x^{T}\Gamma\Sigma_{E}^{-1}\Gamma^{T}x}{x^{T}\Gamma\Gamma^{T}x}\frac{x^{T}\Gamma\Gamma^{T}x}{x^{T}x}\\ &\geq\lambda_{\min}(\Sigma_{E}^{-1})\lambda_{\min}(\Gamma\Gamma^{T})\\ &=\lambda_{\max}(\Sigma_{E})^{-1}\lambda_{\min}(\Gamma)^{2}\end{split}$
Hence,
$\mathbb{E}\left[\frac{1}{n}\|\mathbf{H}\delta-\mathbf{X}b\|_{2}^{2}\right]\leq\|\delta\|_{2}^{2}\lambda_{\max}(\Sigma_{E})\lambda_{\min}(\Gamma)^{-2}=\mathcal{O}\left(\frac{\|\delta\|_{2}^{2}}{p}\right)$
Code:
All the code used for this paper is available here: https://github.com/markusul/SDForest-Paper
R-package for SDForests:
R-package SDModels provides software for non-linear spectrally deconfounded models: https://github.com/markusul/SDModels